OpenStudy (ksaimouli):

2x-1/3x-2<=1 solve the inequalities

OpenStudy (anonymous):

I guess the expression is (2x-1)/(3x-2)<=1. If it is right, so you can write the expression as (-x+1)/(3x-2)<=0. In the upper part the expression is zero if x=1. By other hand x can not be equal to 2/3 (see the bottom part). write now the lines The first one you put the value 1. The second line you put below the first line and put the value 2/3. Of course 2/3 < 1 so in the third line, below the second you put 2/3 and 1. Well int he first line left to 1 is + and right to 1 is - because the expression is (-x+1). The second line, left to 2/3 is - and right to 2/3 is +. Since the expression is (-x+1)/(3x-2) then the sign rules for multiplication is valid to division. Remember + + = + , + - = -, - - = +. Then x< 2/3 is -, 2/3 < x<=1 is + and x>1 is -. So the answer is x<2/3 or x>=1.

OpenStudy (anonymous):

Shouldn't this be in the math section? o_O

OpenStudy (anonymous):

You would have probably gotten more answer if you would of asked it in the math subject section. :D